OpenStudy (anonymous):

What is the period, in seconds, of a simple pendulum of length 2 meters? Use the gravitational constant g = 9.8 m/s2 and round your answer to two decimal places.

OpenStudy (anonymous):

@omnomnom

OpenStudy (omnomnom):

hm... why am i tagged i am of no help to you im so so sorry :(

OpenStudy (anonymous):

aww okay

OpenStudy (omnomnom):

sorry :D

OpenStudy (anonymous):

A physics formula sheet would come in handy right about now . . .

OpenStudy (anonymous):

I think it's sqrt(L/g) though. . .

OpenStudy (anonymous):

The units check out anyway \[\sqrt{\frac{meters}{meters/seconds^2}} = \sqrt{seconds^2} = seconds\]

OpenStudy (anonymous):

here is the equation

OpenStudy (anonymous):

\[d(t) = acos(\omega t) \]

OpenStudy (anonymous):

its the harmonic motion equation

OpenStudy (anonymous):

a period is \[(2\pi)/\omega \]

OpenStudy (anonymous):

Yes, but that is in terms of the amplitude of the swing and the frequency, which are irrelevant.

OpenStudy (anonymous):

The only things that affect the period of a pendulum are its length and the gravitational field strength. You could probably derive the period from the wave equation, but there are too many unknowns.

OpenStudy (anonymous):

I am not like this question lol

OpenStudy (anonymous):

liking* lol

OpenStudy (anonymous):

After finding the period, you could get the frequency and then determine displacement at any time, t for a particular amplitude if you wanted to, but that wasn't your original question. If you just want the period, it's T=sqrt(L/g) like I said earlier.

OpenStudy (anonymous):

i haven't ever used that equation before tbh... can you perhaps help me?

OpenStudy (anonymous):

Are you actually required to derive the period from the harmonic motion equation, or did you just grab that as a possible formula to use?

OpenStudy (anonymous):

thats what equation that goes with particular study so i figured to use that ._.

OpenStudy (anonymous):

Hmm, I could show you how to do it that way (it's a pain in the neck . . .), but first I want to explain the formula I gave.

OpenStudy (anonymous):

Let us avoid pain in thy necks? :3

OpenStudy (anonymous):

Here's how a physicist would think about this problem. The information given is 1. sinmple pendulum 2. length of pendulum (in meters) 3. gravitational acceleration (in meters per second per second) The information asked for is 1. period (in seconds) So, how can we combine the given information (meters, and meters-per-second-squared) to get an answer in seconds?

OpenStudy (anonymous):

with that equation i understand the basics of the equation it self i think ...meters over meters cancels ..and considering seconds is squared the square root will take care of it leaving the answer in seconds right?...its just solving it for seconds

OpenStudy (anonymous):

Exactly. If you didn't happen to have the formula in front of you (I didn't), you'd have to do what I did. I thought about what sequence of arithmetic operations would transform m and m/s^2 into s. And the process is just how you described. One has to divide m by m to cancel that, then square-root the square seconds. The formula then gives itself as sqrt(L/g).

OpenStudy (anonymous):

Plug in your given information and let me know what you get.

OpenStudy (anonymous):

um so sqrt(2/9.8)?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

so that would be .45?

OpenStudy (anonymous):

i tried 2/9.8 and to the root ..and got it wrong i guess i set it up wrong..it was 2. something anyways thnx

OpenStudy (anonymous):

0.45 seconds is the period rounded to 2 places. What do you mean by 2.something? What was 2.something?

OpenStudy (anonymous):

Oh, crap, I forgot to multiply by 2π to convert the units. Sorry about that. Never mind, yes, the period should be 2.84